OpenStudy (darkbluechocobo):

Find the missing value. y is directly proportional to x and inversely proportional to the square of z. y = 20 when x = 50 and z = 5. Find y when x = 3 and z = 6.

OpenStudy (darkbluechocobo):

i know this equation is x=y/z

OpenStudy (campbell_st):

so this is 2 parts directly proportional y = kx substitute y = 20 and x = 50 to find k now the inverse proportion y = k/z^2 to find k, substitute y = 20 and z = 5 they you can complete the rest of the question

OpenStudy (darkbluechocobo):

@campbell_st

OpenStudy (darkbluechocobo):

would this be 20=k/25?

OpenStudy (darkbluechocobo):

(20*25)=k

OpenStudy (campbell_st):

that's correct